# Observation of a non-Hermitian supersonic mode

### v1

@@ Abstract
Quantum computers have long been anticipated to excel in simulating quantum many-body physics. While most previous work has focused on Hermitian physics, we demonstrate the power of variational quantum circuits for resource-efficient simulations of dynamical and equilibrium physics in non-Hermitian systems, revealing new phenomena beyond standard Hermitian quantum machines. Using a variational quantum compilation scheme for fermionic systems, we reduce gate count, save qubits, and eliminate the need for postselection, a major challenge in simulating non-Hermitian dynamics via standard Trotterization. Experimentally, we observed a supersonic mode in the connected density-density correlation function on an $n=18$ fermionic chain after a non-Hermitian, locally interacting quench, which would otherwise be forbidden by the Lieb-Robinson bound in a Hermitian system. Additionally, we investigate sequential quantum circuits generated by tensor networks for ground state preparation, here defined as the eigenstate with the lowest real part eigenvalue, using a variance minimization scheme. Through a trapped-ion implementation on the Quantinuum H1 quantum processor, we accurately capture correlation functions and energies across an exceptional point on a dissipative spin chain up to length $n=20$ using only 3 qubits. Motivated by these advancements, we provide an analytical example demonstrating that simulating single-qubit non-Hermitian dynamics for $\Theta(\log(n))$ time from certain initial states is exponentially hard on a quantum computer, offering insights into the opportunities and limitations of using quantum computation for simulating non-Hermitian physics. Our work raises a number of intriguing questions about the simulability of non-Hermitian systems and the intrinsic properties that enable their efficient quantum simulation, heralding a promising future for quantum computational exploration of such complex quantum phenomena.
@@
The study of non-Hermitian physics traces back to the complex field theory approach by Yang and Lee, who used it to explore new phases of matter [2, 3]. Since then, it has been established that Hamiltonian non-Hermicity leads to unconventional phase transitions and unique phenomena [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], such as exceptional points [18, 19, 20, 21, 22, 23, 24, 25], the non-Hermitian skin effect [26, 27, 28, 29, 30, 31, 32, 33, 34], supersonic modes [35, 36, 37, 38], topological phases [39, 40, 21, 23, 41, 42, 43, 13, 44, 24, 25, 45, 27, 31, 46, 30], and unique entanglement behaviors [47, 48, 49, 34, 50]. Although these discoveries have generated new possibilities and excitement in the era of quantum information [51, 52, 53, 15, 16, 48, 54, 55], non-Hermiticity poses an engineering challenge for conventional electrical, mechanical, photonic and cold atom platforms, and experimental demonstrations have been largely confined to few-particle problems [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 32, 67, 68, 69].
@@
The recent progress of programmable quantum computers is expected to offer new opportunities to solve computational tasks [70], enable cryptography [71], sample hard distributions [72, 73, 74, 75], and, crucially, to simulate many-body physics [76]. However, the presence of physical noise creates a significant obstacle in the faithful execution of such tasks; moreover, the postselection nature of non-Hermitian dynamical simulations adds another layer of complexity. To date, digital quantum simulations of non-Hermitian physics have also been limited in scale [77, 78, 79, 55, 80].
@@
On the other hand, variational quantum algorithms (VQAs) [81] such as the quantum approximate optimization algorithm (QAOA) [82, 83, 84, 85, 86] and the variational quantum eigensolver (VQE) [87, 88, 89, 90, 91, 92, 93, 94] promise to offer advantages in the simulation of quantum systems with near-term quantum hardware [95]. Combined with machine learning [96, 97, 98, 99, 100, 101, 102, 103] and tensor-network techniques [104, 105, 90, 106, 107, 108, 109], VQAs have already yielded successful applications in reducing the resource cost in quantum time evolution [98, 110, 111, 112, 113] and quantum state preparation [90, 114, 115, 116, 108, 109, 117, 118]. It is then natural to ask whether one could utilize these methods in the quantum simulation of non-Hermitian systems.
@@
In this work, we demonstrate the power of variational quantum algorithms through the analytical, numerical, and experimental study of various non-Hermitian systems. First, combining variational quantum compilation (VQC) and Gaussian matrix product state (GMPS) method, we simulate the non-Hermitan quench dynamics in a strongly correlated fermionic system [36] without post-selection, drastically reducing the quantum resources required compared to a Trotterization scheme. On Quantumtinuum’s H1 trapped-ion quantum processor, we observe a non-Hermitian supersonic mode where correlation functions propagate faster than the conventional light cone. Second, by integrating the data compression capabilities of quantum matrix product states (qMPS) [104, 119, 120, 108] with a variance minimization algorithm [121], we devise an eigenstate-finding algorithm for non-Hermitian systems suitable for near-term quantum computers. Testing our algorithm on Quantumtinuum’s H1 trapped-ion quantum computer, we accurately capture the energy and correlation functions of a $n=20$ non-Hermitian spin chain around an exceptional point (EP), where the real-valued ground energy merges with the first excited state to form a Hermitian conjugate pair. This result leverages the qMPS compression capabilities, wherein a classical MPS with bond dimension $\chi$ can be stored on a quantum computer using memory scaling as $q\sim\log\chi$ [104, 108, 117], extending the quantum memory advantages of qMPS to the study of non-Hermitian systems.
@@
Figure 1: (a) A minimal example of a non-Hermitian system: Consider a two-level atom with eigenstates {|$\{\ket{0},\ket{1}\}$. When in the excited state |$\ket{1}$, the atom spontaneously decays to a dark state |$\ket{d}$ at a rate $\gamma$. A cat observer closely monitors the population in the dark state, and by postselecting on the absence of decay, the atom can be effectively thought of as undergoing non-Hermitian time evolution [1, 122]. However, in a direct experimental simulation, the chance of decay not happening (i.e., the probability that postselection successes) would vanish exponentially in time. (b) The use of VQAs in simulating the non-Hermitian quantum quench of Eq. 5 resolves the postselection issue: The half-filled free-fermion ground state is first prepared with a circuit generated by a Gaussian matrix product state. Additional layers of variational unitary circuits are then attached to compress the non-Hermitian time evolution. (c) Numerical observation of a supersonic mode in non-Hermitian quench dynamics. We first find the ground state of Eq. 5 at $J_{z}=0$ with a $\chi=100$ DMRG algorithm and perform a TEBD simulation for the dynamics at $J_{z}=-1.3$. We report the connected density-density correlator between the middle site and the rest of the system. The different modes can be distinguished from the intensity ‘valleys’ of the correlation function. The white solid line denotes the Lieb-Robinson bound and the dashed lines represent the first two supersonic modes; the velocity of the modes can be calculated from a low-energy effective Luttinger theory. Notice that for a longer period of evolution time, the higher energy eigenstates will make the theoretical prediction deviate from the simulation. (d) For the same physical parameter as in (c) at a smaller ($n=18$) system size, we emulate the compressed quench dynamics using a measurement budget of 100,000 shots and report the connected density-density correlator averaged over all bulk sites to enhance signal level. After the GMPS+VQC compression, we find the first supersonic mode remains qualitatively unchanged. (e) Experimental data of the same circuit in (d) collected using Quantinuum’s H1 quantum computer at $t=1$ and $t=2$ for 5,000 shots each. We compare the averaged connected density-density correlator to a classical TEBD simulation. The error bars represent two standard deviations and the shaded blue region indicates the conventional lightcone at $r=2vt$ where $t=2$. The second blue peak concentrated around $r\gg 2vt$ provides clear evidence for information propagating faster than the Lieb-Robinson bound, a unique characteristic to non-Hermitian supersonic physics.
I Observation of a supersonic mode
@@
One of the most striking dynamical properties of non-Hermitian physics is the violation of the Lieb-Robinson (LR) bound [123], a cornerstone of quantum dynamics that imposes strict limits on the propagation of information in many-body systems. Although LR upper bounds the information spreading velocity in a locally interacting Hermitian quantum system, the dynamics under local non-Hermitian Hamiltonians can allow supersonic information propagation modes [35, 36, 124, 125, 126, 50]. This can be probed studying the two-point correlation functions of the quenched dynamics, starting from the ground state of some Hermitian Hamiltonian |$\ket{\psi_{0}}$ and time-evolved with a different non-Hermitian Hamiltonian $H$:
@@
where $N(t)\equiv\bra{\psi(t)}\ket{\psi(t)}$ is the normalization factor, and $\mathcal{O}(t)=e^{iHt}\mathcal{O}e^{-iHt}$ is pseudo-Heisenberg evolution. In (4), the pseudo-Heisenberg evolved operators are restricted by the Lieb-Robinson bound, yet in general the term $e^{iH^{\dagger}t}e^{-iHt}$ can be long-range, enabling supersonic modes even in non-interacting models [36]. In this work, we consider the following non-Hermitian, interacting fermionic chain with open boundary conditions, whose low energy spectrum corresponds to a $\mathcal{PT}$-symmetric Luttinger liquid [36]:
@@
Although $H$ is local, it is possible to show that the system has modes corresponding to the velocity at $x=2kvt$ through a bonsonized effective theory calculation [36], where $k$ is an integer, and
$v=1+\frac{\pi^{2}-8}{8}J_{z}^{2}$
@@
is the LR velocity. This is in strong contrast with the Hermitian case, where only the first ($k=1$) mode is permitted. We study the quench dynamics of the system starting with the free-fermion ground state at $J_{z}=0$ and time evolving the state with a non-zero $J_{z}$ Hamiltonian. We observe the supersonic modes through measuring a connected density-density correlator $C_{nn}^{0,r}(t)=[\langle{n}_{0}{n}_{r}\rangle(t)-\langle{n}_{0}\rangle(t)\langle{n}_{r}\rangle(t)]$, which reveals both the presence of supersonic light cones and their decay, as theoretically predicted and numerically shown in Fig. 1c.
@@
Non-Hermitian time evolution can result from conditional Lindblad-type dynamics, where continuous environmental monitoring is applied to prevent quantum jumps. However, in a direct simulation with, e.g., Trotterizaiton [127, 128] and block encoding [129], the likelihood of post-selection diminishes over time, and thus observing non-Hermitian dynamics becomes increasingly challenging. In fact, the post-selection rate would decay exponentially in $t$ even if one could simulate open system dynamics exactly with an analog simulator. Nevertheless, the suppressed supersonic modes in the correlation function shown in Fig. 1(a) suggest that the post-quench state still exerts some form of locality at least for $t\sim\log(n)$  [^1] For a finite given accuracy, it suffices to truncate the correlation function at a finite mode number and set the cut-off supersonic mode as the effective light cone, which would still be local., providing an opportunity to be compressed by a low complexity state.
@@
In this work, we circumvent this challenge by utilizing two tensor-network-inspired techniques: GMPS [131] for preparing the initial state and VQC [116] for approximating the non-Hermitian time evolution. As we explain in Appx. A, GMPS provides an efficient approximation to the Gaussian mean-field state by exploiting the near-area-law entanglement nature of the ground state, and VQC dramatically compresses time-evolved states that have high entanglement but low complexity. Sketched in Fig. 1(b), we start with a compressed representation of the initial free fermion state and variationally compute a circuit representation at each Trotterized time step, as detailed in Algorithm 1.
Algorithm 1 Compressed quantum quench with VQC
1:Find the GMPS representation that prepares the free-fermion ground state. This could be done either deterministically or variationally [131, 117]. Denote this state as $|\psi_{0}\rangle$.
2:for $t=\Delta t,2\Delta t,...,T$ do
3:     Calculate the normalized target state after a Trotterized time evolution, |$\ket{\psi_{t}}\propto e^{-iHt}\ket{\psi_{t_{0}}}$
4:     Initialize the parametrized quantum circuit.|$\ket{\tilde{\psi}}$ at $\bm{\theta}_{0}$; then use a gradient descent method to minimize infidelity $1-\mathcal{F}=1-|\langle\psi_{t}|\tilde{\psi}(\bm{\theta})\rangle|^{2}$ wrt. $\bm{\theta}$
5:     Record optimized parameter $\bm{\theta}^{*}(t)$, $\bm{\theta}_{0}\leftarrow\bm{\theta}^{*}$
6:     if $1-\mathcal{F}>\epsilon$ then
7:         Add a layer of PQC to the time evolution layers which we initialize at the identity
8:         Return to 4
9:     end if
10:end for
@@
For the parameterized quantum circuit (PQC), we use a parameterized gate known as the ‘fSim’ gate [132], which preserves the $U(1)$ symmetry, as detailed in Appendix C. This choice offers three key benefits: it reduces the number of variational parameters from 15 to 5 per gate, decreases the number of CNOT gates from 3 to 2 per gate, and enables error mitigation with essentially no additional cost. A similar circuit architecture has been proposed in [117] and is denoted as the GMPS+X ansatz. Notice that, generally, the imaginary time evolution does not necessarily conserve charge—only when the initial state is in a single charge sector—which is the situation we consider here. Fixing the target infidelity $0.021-\mathcal{F}=$ 0.02 and comparing our VQC to a naive Trotterization implementation, we find a reduction of CNOT gate count by a factor of 9 and the elimination of auxiliary qubits, reducing their number from 9 to 0. Most significantly, the compressed evolution method does not require postselection, unlike block encoding. As estimated in Appx. C, this avoids a sampling overhead of $10^{19}$.
@@
Another experimental challenge is that supersonic modes have relatively low amplitudes. Increasing $J_{z}$ can increase the intensity of the supersonic modes, but there is a threshold above which the Lutingger Liquid prediction would fail [36]. We set our $J_{z}=-1.3$, right below this threshold. Still, observing the signal level is demanding: In the classical simulation result with time-evolving block decimation (TEBD) Fig. 1 at $n=61$, the correlation strength of the first supersonic model already drops below $\sim 0.01$.
@@
To resolve this, instead of reporting the correlation function $C_{nn}^{i,i+r}(t)$ for a specific initial position in the chain, we report the correlation function averaged over all bulk sites $\overline{C_{nn}^{i,i+r}(t)}$ because the supersonic mode should be a universal property within the system. Asymptotically, this post-processing technique effectively increased the amount of data by a factor of $\sim(n-r)$, roughly an order of magnitude in this case. We illustrate our experimental details in the Appx. C.
@@
It is worth noticing that for the time evolution generated by a non-Hermitian Hamiltonian, completely circumventing postselection with $\operatorname{poly}(n)$ gates is nearly impossible even with the use of variational compilation unless $\mathsf{BQP=PostBQP=PP}$ [133]. More surprisingly, for certain initial states, we find an example where the time evolution of a Hamiltonian consisting only of single qubit terms could be too powerful to simulate, even for a very short period:
Theorem I.1 (Single qubit non-Hermitian dynamics for logarithmic time can be hard to simulate with a quantum circuit).
@@
One example is to take |$\ket{\psi_{0}}$ to be a Haar random state and $H=-i\sum_{i}Z_{i}$: running the time evolution for $t=\Theta(\log(n))$ allows one to distinguish the Haar random state to a maximally mixed state, which should be exponentially hard even for quantum circuits [134]. We provide a detailed analytical proof of this no-go theorem in Appx. A, where we then show that the theorem can be thought of as an oracle separation between $\mathsf{BQP}$ and $\mathsf{PostBQP}$.
II Crossing the exceptional point

### v2

@@ Abstract
Quantum computers have long been anticipated to excel in simulating quantum many-body physics. In this work, we demonstrate the power of variational quantum circuits for resource-efficient simulations of dynamical and equilibrium physics in non-Hermitian systems. Using a variational quantum compilation scheme for fermionic systems, we reduce gate count, save qubits, and eliminate the need for postselection, a major challenge in simulating non-Hermitian dynamics via standard Trotterization. On the Quantinuum H1 trapped-ion processor, we experimentally observed a supersonic mode on an $n=18$ fermionic chain after a non-Hermitian, nearest-neighbor interacting quench, which would otherwise be forbidden in a Hermitian system. Additionally, we investigate sequential quantum circuits generated by tensor networks for ground-state preparation using a variance minimization scheme, accurately capturing correlation functions and energies across an exceptional point on a dissipative spin chain up to length $n=20$ using only 3 qubits. On the other hand, we provide an analytical example demonstrating that simulating single-qubit non-Hermitian dynamics for $\Theta(\log(n))$ time from certain initial states is exponentially hard on a quantum computer. Our work raises many intriguing questions about the intrinsic properties of non-Hermitian systems that permit efficient quantum simulation.
I Introduction
@@
The study of non-Hermitian physics traces back to the complex field theory approach by Yang and Lee, who used it to explore new phases of matter [2]. Since then, it has been established that Hamiltonian non-Hermicity leads to unconventional phase transitions and unique phenomena [3, 4, 5, 6], such as exceptional points [7], the non-Hermitian skin effect [8], supersonic modes [9, 10], topological phases [11], and unique entanglement behaviors [12, 13]. Although these discoveries have generated new possibilities and excitement in the era of quantum information, non-Hermiticity poses an engineering challenge for conventional electrical, mechanical, photonic and cold atom platforms, and experimental demonstrations have been largely confined to few-particle problems [14, 15, 16].
@@
The recent progress of programmable quantum computers is expected to offer new opportunities to solve computational tasks [17], enable cryptography [18], sample hard distributions [19], and, crucially, to simulate many-body physics [20]. However, the presence of physical noise creates a significant obstacle in the faithful execution of such tasks. More essentially, the postselection nature of non-Hermitian dynamical simulations brings the challenge to another level: In a “direct” simulation by Trotterization, the success rate of postselection would vanish exponentially in time, creating a huge sampling overhead. To date, digital quantum simulations of non-Hermitian physics have also been limited in scale compared to those of Hermitian quantum systems [21, 22].
@@
On the other hand, variational quantum algorithms (VQAs) [23] such as the quantum approximate optimization algorithm (QAOA) [24] and the variational quantum eigensolver (VQE) [25] promise to offer advantages in the simulation of quantum systems with near-term quantum hardware [26]. Combined with machine learning [27] and tensor-network techniques [28], VQAs have already yielded successful applications in reducing the resource cost in quantum time evolution [29, 30] and quantum state preparation [31]. It is then natural to ask whether one could utilize these methods in the quantum simulation of non-Hermitian systems. Intuitively, VQAs explore the large degree of freedom in the Hilbert space and could mitigate the postselection issue, and even reduce the circuit gate count.
@@
In this work, we demonstrate the power of variational quantum algorithms through the analytical, numerical, and experimental study of various non-Hermitian systems. First, combining variational quantum compilation (VQC) and Gaussian matrix product state (GMPS) method, we simulate the non-Hermitan quench dynamics in a strongly correlated fermionic system [10] without post-selection, drastically reducing the quantum resources required compared to a Trotterization scheme. Most significantly, this includes a $\sim 10^{19}$ sampling overhead due to post-selection, as we estimate. On Quantumtinuum’s H1 trapped-ion quantum processor, we observe a non-Hermitian supersonic mode where correlation functions propagate faster than the conventional light cone. Second, by integrating the data compression capabilities of quantum matrix product states (qMPS) [28, 32] with a variance minimization algorithm, we devise an eigenstate-finding algorithm for non-Hermitian systems suitable for near-term quantum computers. Testing our algorithm on Quantumtinuum’s H1 trapped-ion quantum computer, we accurately capture the energy and correlation functions of a $n=20$ non-Hermitian spin chain around an exceptional point (EP), where the real-valued ground energy merges with the first excited state to form a Hermitian conjugate pair. This result leverages the qMPS compression capabilities, wherein a classical MPS with bond dimension $\chi$ can be stored on a quantum computer using memory scaling as $q\sim\log\chi$ [28, 31], extending the quantum memory advantages of qMPS to the study of non-Hermitian systems.
@@
Figure 1: (a) A minimal example of a non-Hermitian system: Consider a two-level atom with eigenstates {|$\{\ket{0},\ket{1}\}$. When in the excited state |$\ket{1}$, the atom spontaneously decays to a dark state |$\ket{d}$ at a rate $\gamma$. A cat observer closely monitors the population in the dark state, and by postselecting on the absence of decay, the atom can be effectively thought of as undergoing non-Hermitian time evolution [1]. (b) The use of VQAs in simulating the non-Hermitian quantum quench of Eq. 5 resolves the postselection issue: The half-filled free-fermion ground state is first prepared with a circuit generated by a Gaussian matrix product state. Additional layers of variational unitary circuits are then attached to compress the non-Hermitian time evolution. (c) Numerical observation of a supersonic mode in non-Hermitian quench dynamics. The different modes can be distinguished from the intensity ‘valleys’ of the correlation function. The white solid line denotes the Lieb-Robinson bound and the dashed lines represent the first two supersonic modes. (d) For the same physical parameter as in (c) at $n=18$, we emulate the compressed quench dynamics using 100,000 shots and report the connected density-density correlator averaged over bulk sites to enhance signal level. (e) Experimental data of the same circuit in (d) collected using Quantinuum’s H1 quantum computer at $t=1$ and $t=2$ for 5,000 shots each. We compare the averaged connected density-density correlator to a classical TEBD simulation. The error bars represent two standard deviations and the shaded blue region indicates the conventional lightcone at $r=2vt$ where $t=2$.
II Results
II.1 Observation of a supersonic mode
@@
One of the most striking dynamical properties of non-Hermitian physics is the violation of the Lieb-Robinson (LR) bound [33], a cornerstone of quantum dynamics that imposes strict limits on the propagation of information in many-body systems. Although LR upper bounds the information spreading velocity in a locally interacting Hermitian quantum system, the dynamics under strictly geometrically two local non-Hermitian Hamiltonians can allow supersonic information propagation modes [9], without introducing any (quasi) long-range interactions [34, 35]. This can be probed by studying the two-point correlation functions of the quenched dynamics, starting from the ground state of some Hermitian Hamiltonian |$\ket{\psi_{0}}$ and time-evolved with a different non-Hermitian Hamiltonian $H$:
@@
where $N(t)\equiv\bra{\psi(t)}\ket{\psi(t)}$ is the normalization factor, and $\mathcal{O}(t)=e^{iHt}\mathcal{O}e^{-iHt}$ is pseudo-Heisenberg evolution. In (4), the pseudo-Heisenberg evolved operators are restricted by the Lieb-Robinson bound, yet in general the term $e^{iH^{\dagger}t}e^{-iHt}$ can be long-range, enabling supersonic modes even in non-interacting models. In this work, we consider the following non-Hermitian, interacting fermionic chain with open boundary conditions, whose low energy spectrum corresponds to a $\mathcal{PT}$-symmetric Luttinger liquid [10, 36]:
@@
Although $H$ is local, it is possible to show that the system has modes corresponding to the velocity at $x=2kvt$ through a bosonized effective theory calculation [10], where $k$ is an integer, and
$v=1+\frac{\pi^{2}-8}{8}J_{z}^{2}$ (6)
@@
is the LR velocity. This is in strong contrast with the Hermitian case or systems in a full Linbladian evolution [9], where only the first ($k=1$) mode is permitted. We study the quench dynamics of the system starting with the free-fermion ground state at $J_{z}=0$ and time evolving the state with a non-zero $J_{z}$ Hamiltonian. We observe the supersonic modes through measuring a connected density-density correlator $C_{nn}^{0,r}(t)=[\langle{n}_{0}{n}_{r}\rangle(t)-\langle{n}_{0}\rangle(t)\langle{n}_{r}\rangle(t)]$, which reveals both the presence of supersonic light cones and their decay, as theoretically predicted and numerically shown in Fig. 1c.
@@
Non-Hermitian time evolution can result from conditional Lindblad-type dynamics, where continuous environmental monitoring is applied to prevent quantum jumps. However, in a direct simulation with, e.g., Trotterizaiton [37] and block encoding [38], the likelihood of post-selection diminishes over time, and thus observing non-Hermitian dynamics becomes increasingly challenging. In fact, the post-selection rate would decay exponentially in $t$ even if one could simulate open system dynamics exactly with an analog simulator. Nevertheless, the suppressed supersonic modes in the correlation function shown in Fig. 1(a) suggest that the post-quench state still exerts some form of locality at least for $t\sim\log(n)$, providing an opportunity to be compressed by a low complexity state. For a finite given accuracy, it suffices to truncate the correlation function at a finite mode number and set the cut-off supersonic mode as the effective light cone, which would still be efficiently simulable.
@@
In this work, we circumvent this challenge by utilizing two tensor-network-inspired techniques: GMPS [39] for preparing the initial state and VQC [29] for approximating the non-Hermitian time evolution. As we explain in the supplement, GMPS provides an efficient approximation to the Gaussian mean-field state by exploiting the near-area-law entanglement nature of the ground state, and VQC dramatically compresses time-evolved states that have high entanglement but low complexity. Sketched in Fig. 1(b), we start with a compressed representation of the initial free fermion state and variationally compute a circuit representation at each Trotterized time step, as detailed in Algorithm 1. In practice, we found that it suffices to use a shallow, unitary variational circuit to capture the action of the non-Hermitian dynamics, without the necessity to introduce ancillary qubits.
Box 1 – Compressed quantum quench with VQC 1:Find the GMPS representation that prepares the free-fermion ground state. This could be done either deterministically or variationally [39, 40]. Denote this state as $|\psi_{0}\rangle$. 2:for $t=\Delta t,2\Delta t,...,T$ do 3:     Calculate the normalized target state after a Trotterized time evolution, |$\ket{\psi_{t}}\propto e^{-iHt}\ket{\psi_{t_{0}}}$ 4:     Initialize the parametrized quantum circuit.|$\ket{\tilde{\psi}}$ at $\bm{\theta}_{0}$; then use a gradient descent method to minimize infidelity $1-\mathcal{F}=1-|\langle\psi_{t}|\tilde{\psi}(\bm{\theta})\rangle|^{2}$ wrt. $\bm{\theta}$ 5:     Record optimized parameter $\bm{\theta}^{*}(t)$, $\bm{\theta}_{0}\leftarrow\bm{\theta}^{*}$ 6:     if $1-\mathcal{F}>\epsilon$ then 7:         Add a layer of PQC to the time evolution layers which we initialize at the identity 8:         Return to 4 9:     end if 10:end for
@@
For the parameterized quantum circuit (PQC), we use a parameterized gate known as the ‘fSim’ gate [41], which preserves the $U(1)$ symmetry, as detailed in Supplementary Note 1. This choice offers three key benefits: it reduces the number of variational parameters from 15 to 5 per gate, decreases the number of CNOT gates from 3 to 2 per gate, and enables error mitigation with essentially no additional cost. A similar circuit architecture has been proposed in [40] and is denoted as the GMPS+X ansatz. Notice that, generally, the imaginary time evolution does not necessarily conserve charge—only when the initial state is in a single charge sector—which is the situation we consider here. Fixing the target infidelity $0.021-\mathcal{F}=$ 0.02 and comparing our VQC to a naive Trotterization implementation, we find a reduction of CNOT gate count by a factor of 9 and the elimination of auxiliary qubits, reducing their number from 9 to 0. Most significantly, the compressed evolution method does not require postselection, unlike block encoding. As estimated in the Supplementary Note 1, this avoids a sampling overhead of $10^{19}$.
@@
Another experimental challenge is that supersonic modes have relatively low amplitudes. Increasing the magnitude of $J_{z}$ can increase the intensity of the supersonic modes, but there is a threshold $\sim-1.33$, above which the Lutingger Liquid prediction would fail [10] and thus we set our $J_{z}=-1.3$. Still, observing the signal level is demanding: In the classical simulation result with time-evolving block decimation (TEBD) Fig. 1 at $n=61$, the correlation strength of the first supersonic model already drops below $\sim 0.01$.
@@
To resolve this, instead of reporting the correlation function $C_{nn}^{i,i+r}(t)$ for a specific initial position in the chain, we report the correlation function averaged over all bulk sites $\overline{C_{nn}^{i,i+r}(t)}$ because the supersonic mode should be a universal property within the system. Asymptotically, this post-processing technique effectively increased the amount of data by a factor of $\sim(n-r)$, roughly an order of magnitude in this case. We illustrate our experimental details in Supplementary Note 1.
@@
It is worth noticing that for the time evolution generated by a generic non-Hermitian Hamiltonian, completely circumventing postselection with $\operatorname{poly}(n)$ gates is nearly impossible even with the use of variational compilation unless complexity classes collapses: $\mathsf{BQP=PostBQP=PP}$ [42]. But what if we make strong assumptions about the Hamiltonians, such as they are physically local? Surprisingly, for certain initial states, we find an example where the time evolution of a Hamiltonian consisting only of single qubit terms could be too powerful to simulate, even for a very short period:
Theorem II.1 (Single qubit non-Hermitian dynamics for logarithmic time can be hard to simulate with a quantum circuit).
@@
One example is to take |$\ket{\psi_{0}}$ to be a Haar random state and $H=-i\sum_{i}Z_{i}$: running the time evolution for $t=\Theta(\log(n))$ allows one to distinguish the Haar random state to a maximally mixed state, which should be exponentially hard even for quantum circuits [43]. We provide a detailed analytical proof of this no-go theorem in the Methods section. As a corollary, we also show that the theorem can be thought of as an oracle separation between complexity classes $\mathsf{BQP}$ and $\mathsf{PostBQP}$, as well as between $\mathsf{BQP}$ and $\mathsf{PostQNC_{0}}$ (to be defined in the Methods).
II.2 Crossing the exceptional point